OpenStudy (anonymous):

For questions 3 and 4, write a rule for the sequence. 3) 1, 2, 4, 8, . . . A)Start with 1 and then add 1 repeatedly. B)Start with 1 and then add multiples of 2 repeatedly. C)Start with 1 and then multiply by 2 repeatedly. D)Start with 1 and then divide by 2 repeatedly. 4) 50, 40, 30, 20, . . . A)Start with 50 and then add 10 repeatedly. B)Start with 50 and then add −10 repeatedly. C)Start with 20 and then add 10 repeatedly. D)Start with 50 and then add multiples of 10 repeatedly.

OpenStudy (anonymous):

HELPPP

OpenStudy (anonymous):

first one looks like C right? \(1, 1\times 2, 2\times 2, 2\times 2\times 2...\)

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

second one looks like you are subtracting ten each time, which some weirdo math teacher wrote as "adding minus ten"

OpenStudy (anonymous):

LOL OK